# Graph-Theoretic Analysis of $n$-Replica Time Evolution in the Brownian Gaussian Unitary Ensemble

### v1

@@ 4 Time Evolution for $n=3$
Then we consider $n=3$, so we have six sites: $\{1,\bar{1},2,\bar{2},3,\bar{3}\}$. The permutation group is $S_{3}\otimes S_{3}$, $S_{3}=\{I,X_{12},X_{13},X_{23},X_{123},X_{132}\}$, then
@@ 4.1 Grouping Graphs into Categories
Here we first define some notations:
@@
$N_{3,p}=\{36,324,324,36\}\ ,\quad\text{for }p=\{0,1,2,3\}\ \ .$ (64)
Next, we discuss the graph categories $\mathsf{F}_{p,\alpha}$ with $p$ connected pairs. For a graph with $p$ pairs, notice that the second term in (23) ($\sum_{u\cap C[\mathsf{F}]=\emptyset}c_{u}\mathsf{F}$) represents graphs with $p+1$ pairs, while the last two terms indicate applying additional 2 swaps on $\mathsf{F}$, which are excluded by $C(\mathsf{F})$.
@@
  • For $p=1$, with $C[\mathsf{F}]=c_{u}$, we should require $X\cap u=\emptyset$. We can define the allowed set $X^{(0)}=X^{(ex)}[u]\equiv\{X|X\cap u\neq\emptyset\}$ and its complement as $X^{(1)}$. For example, if $u=(1\bar{1})$, then we have $X^{(0)}=\{X_{23}\}$ and $X^{(1)}=X-X^{(0)}=\{X_{12},X_{13}\}$.
@@
While the discussion of graph categories by connected pairs is complete for $n=3$, we could generalize the definitions of $X^{(0)}$ and $X^{(1)}$ to $X^{(\alpha)}$ for generic $n$ and $p$. We define the conditions set as follows
@@
$\text{Con}(0,A,u)\equiv(A\cap u=\emptyset),~{}\text{Con}(1,A,u)\equiv(A\cap u\not=\emptyset)\ ,$ (66)
showing that whether $u$ intersects with $A$ or not. Hence for a series of certain connected pairs $u_{1},u_{2},\ldots,u_{p}$, $X^{(\alpha)}$ is a set of 2-swap operators:
@@
In our case, for instance, we take $n=3,p=2$, then the condition set is $\textsf{Cons}=\{00,01,10,11\}$. And if we further choose $u_{1}=(1\bar{1}),u_{2}=(2\bar{2})$, we have
@@
Equipped with the new categories $\mathsf{F}_{0,I}$, we could calculate the operations of $\mathcal{L}_{n}^{r}$ on those categories. We will find that $\mathcal{L}_{n}^{r}\mathsf{F}_{0,I}$ can be expressed as a linear combination of graphs represented by permutations
@@
where the product should be understood as the product between sets, which means the repeated terms should be eliminated. And the summation runs over all possible permutations. We will primarily discuss how to assign these blocks (69) into the appropriate graph categories in the followings.
Zero Pair: $C[\mathsf{F}]=0$, #$\#\mathsf{F}=36$
For $p=0$, (69) becomes
@@
Notice that for $n=3$, we have the following relation for permutation sets
$XX=3I+3[XX],~{}[XX]X=2X\ .$ (71)
Hence all terms appearing in $(X+\bar{X})^{m}$ are
@@
Since the gauge group vanishes, there is no redundancy. Hence we obtain $6$ categories
$\mathsf{F}_{0,I},\mathsf{F}_{0,X+\bar{X}},\mathsf{F}_{0,X\bar{X}},\mathsf{F}_{0,[XX]+[\bar{X}\bar{X}]},\mathsf{F}_{0,[XX]\bar{X}+[\bar{X}\bar{X}]X},\mathsf{F}_{0,[XX][\bar{X}\bar{X}]}\ ,$ (73)
@@
$\mathsf{F}_{0,I}\equiv c[I],\quad\mathsf{F}_{0,X+\bar{X}}\equiv\sum_{e}c[X_{e}]+\sum_{\bar{e}}c[X_{\bar{e}}]+(X\leftrightarrow\bar{X})\ ,$ (74)
$\mathsf{F}_{0,X\bar{X}}\equiv\sum_{u,\bar{e}}c[X_{\bar{e}}X_{u}]+(X\leftrightarrow\bar{X})\ ,$
$\mathsf{F}_{0,[XX]+[\bar{X}\bar{X}]}\equiv c[X_{123}]+c[X_{132}]+c[X_{\bar{1}\bar{2}\bar{3}}]+c[X_{\bar{1}\bar{3}\bar{2}}]+(X\leftrightarrow\bar{X})\ ,$
$\mathsf{F}_{0,[XX]\bar{X}+[\bar{X}\bar{X}]X}\equiv\sum_{e}c\left[X_{\bar{1}\bar{2}\bar{3}}X_{e}+X_{\bar{1}\bar{3}\bar{2}}X_{e}\right]+\sum_{\bar{e}}c\left[X_{123}X_{\bar{e}}+X_{132}X_{\bar{e}}\right]+(X\leftrightarrow\bar{X})\ ,$
$\mathsf{F}_{0,[XX][\bar{X}\bar{X}]}\equiv c\left[X_{123}X_{\bar{1}\bar{2}\bar{3}}\right]+c\left[X_{123}X_{\bar{1}\bar{3}\bar{2}}\right]+c\left[X_{132}X_{\bar{1}\bar{2}\bar{3}}\right]+c\left[X_{132}X_{\bar{1}\bar{3}\bar{2}}\right]+(X\leftrightarrow\bar{X})\ .$
It is straightforward to count their sizes, which are $\{1,6,9,4,12,4\}$ respectively. Thus, we have $1+6+9+4+12+4=36$ graphs, as we have calculated in (64).
One Pair: $C[\mathsf{F}]=1$, #$\#\mathsf{F}=324$
For $p=1$, (69) becomes
@@
The size of the gauge group is just $p!=1$, so there is no redundancy. For arbitrary $u=(i\bar{j})$, we find that
#$\#X^{(0)}=1,~{}\#X^{(1)}=2,~{}X=X^{(0)}+X^{(1)}\ .$ (76)
Hence we have
@@
with there sizes to be $\{1,18,36,9,36,36,36,36,72,36\}$.
Two Pairs: $C[\mathsf{F}]=2$,#$\#\mathsf{F}=324$
For $p=2$, (69) becomes
@@
$\forall u,v,\ u\cap v=\emptyset$, we have
@@
$c_{u}c_{v}c[\sigma]=c_{u}c_{v}c\left[\sigma\cdot X^{(11)}\bar{X}^{(11)}\right]\ ,$ (83)
@@
Their sizes are $\{18,72,18,36,36,72,72\}$.
Three Pairs: $C[\mathsf{F}]=3$,#$\#\mathsf{F}=36$
For $p=2$, (69) becomes
@@
For three pairs $u_{1},u_{2},u_{3}$, we can only take three conditions
@@
Any permutation $X_{e}X_{\bar{e}}$ is considered as a redundancy, allowing us to fix the order of $\bar{1},\bar{2},\bar{3}$. This means we only need to consider permutations of $1,2,3$. Utilizing the fact that $X=X^{(011)}+X^{(101)}+X^{(110)}$, we identify three building blocks: $I,X,[XX]$. The respective graph categories are:
$\mathsf{F}_{3,I}=\cup_{u_{1},u_{2},u_{3}}c[I],~{}\mathsf{F}_{3,X}=\cup_{u}c_{u_{1}}c_{u_{2}}c_{u_{3}}c[X],~{}\mathsf{F}_{3,XX}=\cup_{u_{1},u_{2},u_{3}}c_{u_{1}}c_{u_{2}}c_{u_{3}}c[[XX]]\ .$ (88)
They have sizes of $\{6,18,12\}$, respectively.
4.2 Operation of $\mathcal{L}_{3}$
@@
$XX_{e}$ $=X_{e}+[XX],~{}[XX]X_{e}=X_{e}[XX]=X-X_{e}\ ,$
and the gauge symmetry discussed earlier, it is straightforward to calculate the operation of $\mathcal{L}_{3}$ on the graph categories defined in the previous section. For simplicity, we define $\mathcal{L}_{n}^{\prime}$ as the non-trivial part of $\mathcal{L}_{n}$:
@@
$\mathcal{L}_{3}\mathsf{F}_{3,\beta}=(w+3J)\mathsf{F}_{3,\beta},\quad\beta=I,X,XX\ .$ (94)
@@ 4.3 Observables
Recall the factor $\mathsf{w}=-\text{i}\mathsf{E}-nJ$ from (13), then we have $M=-\text{i}\mathsf{E}\mathbf{1}+M_{J}$. Here $\mathbf{1}$ denotes identity matrix. $M_{J}$ has the explicit form
@@
It is straightforward to obtain the spectrum of $M_{J}$:
[…]
@@
Then we can find the analytical expression of $\mathcal{U}_{3}(t)$ as
$\mathcal{U}_{3}(t)=e^{\mathcal{L}_{3}t}=e^{-\text{i}\mathsf{E}t}\sum_{a=1}^{26}f_{a}(t)\mathsf{F}_{a}\ .$ (123)
@@
Figure 3: The pattern coefficient $f_{a}(t)$ and its log-plot, where we set $D=10$ and $J=1$. We depict the categories of $p=0,1,2,3$ with red, orange, green, and blue colors, respectively. One can see that the three categories with $p=3$ have energy 00, so they (the blue lines) remain finite at time $t=\infty$.
Generally, we can calculate the correlation functions as
$\text{Tr}\left[O_{1}(t)O_{2}O_{3}(t)O_{4}O_{5}(t)O_{6}\right]=\leavevmode\hbox{$\vbox{\hbox{\resizebox{}{}{{\leavevmode\hbox{\set@color{}}}}}}$}\equiv\sum_{a=1}^{26}f_{a}(t)\times\left(e^{-\text{i}\mathsf{E}t}\mathsf{F}_{a}\right)_{O_{1},O_{2},O_{3},O_{4},O_{5},O_{6}}$ (124)
where we have used $O(t)\equiv U^{\dagger}OU$ and the graph representation for an operator $A$
$A_{ij}=\leavevmode\hbox{$\vbox{\hbox{\resizebox{}{}{{\leavevmode\hbox{\set@color{}}}}}}$}\ .$ (125)
To calculate $\mathsf{F}_{a}[O_{1},O_{2},O_{3},O_{4},O_{5},O_{6}]$, in principle, we need to contract each individual graph in the category with the operator tensor in the way as shown in (124). The expression $\mathsf{F}=\prod_{i=1}^{p}c_{a_{i}\bar{b}_{i}}c[\sigma]$ has told us the connection between points in the graph. We treat the connections in the graph as free propagators and the operators as bridges between points $i,\bar{i}$ and $\bar{i}^{\prime},\overline{(i+1)}^{\prime}$, where the prime means that the point is on the right-hand side of the graph. So there is no need to draw the graph. We can just start with any points following the connections in the graph and form a closed loop. Hence we will have a trace of operators
@@
If there are still points that we have not passed, we then start with those points, finding another closed loop until all points in the graph have been visited. The process may be a bit tedious, but it can be accomplished with some code. Alternatively, one can simply write down the explicit expression of a graph using the Kronecker delta function as defined in (16).
@@
$\mathbb{E}\left(\mathbb{I}_{D}^{\otimes 2}\otimes\mathbb{I}_{D}^{\otimes 2}\otimes U(t_{32})\otimes U^{*}(t_{32})\right)\ \mathbb{E}\left(\mathbb{I}_{D}\otimes\mathbb{I}_{D}\otimes U(t_{21})^{\otimes 2}\otimes U^{*}(t_{21})^{\otimes 2}\right)\ \mathbb{E}\left(U(t_{1})^{\otimes 3}\otimes U^{*}(t_{1})^{\otimes 3}\right)$
$=$ $\left[\mathbb{I}_{D}^{\otimes 2}\otimes\mathbb{I}_{D}^{\otimes 2}\otimes\mathcal{U}_{1}(t_{32})\right]\times\left[\mathbb{I}_{D}\otimes\mathbb{I}_{D}\otimes\mathcal{U}_{2}(t_{21})\right]\times\mathcal{U}_{3}(t_{1})\ ,$ (127)
where $\mathbb{I}_{D}$ denotes the $D$-dimensional identity matrix. Finally, we simply need to contract $\mathcal{U}_{3}(t_{1},t_{2},t_{3})$ with the operator tensor as demonstrated in (124). In this paper, we do not provide explicit expressions for the correlation functions.
5 General Cases
For general $n$, we have $2n$ sites $\{1,\bar{1},2,\bar{2},3,\bar{3},\ldots,n,\bar{n}\}$. The permutation group is $S_{n}\otimes S_{n}$. As discussed in $n=3$ case, we should define

### v2

@@ 4 Time evolution for $n=3$
Then we consider $n=3$, so we have six contours: $\{1,\bar{1},2,\bar{2},3,\bar{3}\}$. The permutation group is $S_{3}\otimes S_{3}$, $S_{3}=\{I,X_{12},X_{13},X_{23},X_{123},X_{132}\}$, then
@@ 4.1 Grouping graphs into categories
Here we first define some denotations:
@@
$N_{3,p}=\{36,324,324,36\}\ .$ (58)
Next, we discuss the categories $\mathsf{F}_{p,\alpha}$ for $p=\{0,1,2,3\}$. For a graph with $p$ pairs, observe that the second term in Eq.(22) ($\sum_{u\cap C[\mathsf{F}]=\emptyset}c_{u}\mathsf{F}$) represents graphs with $p+1$ pairs, while the last two terms indicate applying additional 2-cycles, which are excluded by $C(\mathsf{F})$.
@@
  • For $p=1$, with $C[\mathsf{F}]=c_{u}$, we require $X\cap u=\emptyset$. We denote this as $X^{(ex)}[u]$, or simply $X^{(0)}$. Similarly, we can define $X^{(0)}=X^{(ex)}[u]\equiv\{X|X\cap u\neq\emptyset\}$. For example, if $u=(1\bar{1})$, then $X^{(0)}=\{X_{23}\}$ and $X^{(1)}=X-X^{(0)}=\{X_{12},X_{13}\}$.
@@
We can generalize the definitions of $X^{(0)}$ and $X^{(1)}$ for a general $p$. We define the conditions set as follows
@@
[…]
and for a certain connected pair $c(u_{1},u_{2},\ldots,u_{p})$, $X^{(\alpha)}$ is a set of two-cycles:
@@
Here we take $n=3,p=2$ and $u_{1}=(1\bar{1}),u_{2}=(2\bar{2})$, then the condition set is $\textsf{Cons}=\{00,01,10,11\}$, so
@@
Starting with $\mathsf{F}_{0,I}$, we will find that $\mathcal{L}_{n}^{r}\mathsf{F}_{0,I}$ can be expressed as a linear combination of
@@
where the product in square brackets should be understood as the product between permutation sets, which we then sum over all permutations. The condition indices $\alpha_{j}$ belong to the conditions set. Next, we will primarily discuss how to classify these blocks (Eq.(63)) into the appropriate graph categories.
Zero pair: $C[\mathsf{F}]=0$, $\mathsf{F}=36$
For $p=0$, Eq.(63) becomes
@@
Notice that there is a relation for $n=3$
$XX=3I+3[XX],~[XX]X=2X\ ,$ (65)
so all terms appearing in $(X+\bar{X})^{m}$ are
@@
Since the gauge group vanish here, there is no redundancy, then we obtain $6$ categories
$\mathsf{F}_{0,I},\mathsf{F}_{0,X+\bar{X}},\mathsf{F}_{0,X\bar{X}},\mathsf{F}_{0,[XX]+[\bar{X}\bar{X}]},\mathsf{F}_{0,[XX]\bar{X}+[\bar{X}\bar{X}]X},\mathsf{F}_{0,[XX][\bar{X}\bar{X}]}$ (67)
@@
$\mathsf{F}_{0,I}$ $\equiv c[I],\mathsf{F}_{0,X}\equiv\sum_{e}c[X_{e}]+\sum_{\bar{e}}c[X_{\bar{e}}]\ ,$ (68)
$\mathsf{F}_{0,X\bar{X}}$ $\equiv\sum_{u,\bar{e}}c[X_{\bar{e}}X_{u}],\mathsf{F}_{0,XX}\equiv c[X_{123}]+c[X_{132}]+c[X_{\bar{1}\bar{2}\bar{3}}]+c[X_{\bar{1}\bar{3}\bar{2}}]\ ,$
$\mathsf{F}_{0,XX\bar{X}}$ $\equiv\sum_{e}c\left[X_{\bar{1}\bar{2}\bar{3}}X_{e}+X_{\bar{1}\bar{3}\bar{2}}X_{e}\right]+\sum_{\bar{e}}c\left[X_{123}X_{\bar{e}}+X_{132}X_{\bar{e}}\right]\ ,$
$\mathsf{F}_{0,XX\bar{X}\bar{X}}$ $\equiv c\left[X_{123}X_{\bar{1}\bar{2}\bar{3}}\right]+c\left[X_{123}X_{\bar{1}\bar{3}\bar{2}}\right]+c\left[X_{132}X_{\bar{1}\bar{2}\bar{3}}\right]+c\left[X_{132}X_{\bar{1}\bar{3}\bar{2}}\right]\ .$
Without introducing complexities, we neglect the dual part obtained by $X\leftrightarrow\bar{X}$ and simply denote the categories containing $[XX]$ and $\bar{X}\bar{X}]$ as $\mathsf{F}_{0,XX}$, $\mathsf{F}_{0,XX\bar{X}}$, and $\mathsf{F}_{0,XX\bar{X}\bar{X}}$. It is straightforward to count their sizes: $\{1,6,9,4,12,4\}$. Thus, we have $1+6+9+4+12+4=36$, as required.
One Pair: $C[\mathsf{F}]=1$$\mathsf{F}=324$
For $p=1$, Eq.(63) becomes
@@
The size of gauge group is just $p!=1$, so there is no redundancy. For arbitrary $u=(i\bar{j})$, we find
$\#X^{(0)}=1,~\#X^{(1)}=2,~X=X^{(0)}+X^{(1)}\ ,$ (70)
so we have
@@
with sizes $\{1,18,36,9,36,36,36,36,72,36\}$.
Two pairs: $C[\mathsf{F}]=2$$\mathsf{F}=324$
For $p=2$, Eq.(63) becomes
@@
For any two pairs $u\cap v=\emptyset$, we have
@@
$c_{u}c_{v}c[\sigma]=c_{u}c_{v}c\left[\sigma\cdot X^{(11)}\bar{X}^{(11)}\right]\ .$ (77)
@@
Their size are $\{18,72,18,36,36,72,72\}$.
Three pairs: $C[\mathsf{F}]=3$$\mathsf{F}=36$
For $p=2$, Eq.(63) becomes
@@
For three pairs $u_{1},u_{2},u_{2}$, we can only take three conditions
@@
Any permutation $X_{e}X_{\bar{e}}$ is considered a redundancy, allowing us to fix the order of $\bar{1},\bar{2},\bar{3}$. This means we only need to consider permutations of $1,2,3$. Utilizing the fact that $X=X^{(011)}+X^{(101)}+X^{(110)}$, we identify the three building blocks: $I,X,[XX]$. The respective graph categories are:
$\mathsf{F}_{3,I}=\cup_{u_{1},u_{2},u_{3}}c[I],\mathsf{F}_{3,X}=\cup_{u}c_{u_{1}}c_{u_{2}}c_{u_{3}}c[X],\mathsf{F}_{3,XX}=\cup_{u_{1},u_{2},u_{3}}c_{u_{1}}c_{u_{2}}c_{u_{3}}c[[XX]]\ .$ (82)
They have sizes $\{6,18,12\}$.
4.2 Action of $\mathcal{L}_{3}$
@@
$XX_{e}$ $=X_{e}+[XX],~[XX]X_{e}=X_{e}[XX]=X-X_{e}\ .$
and the gauge symmetry discussed earlier, it is straightforward to calculate the action of $\mathcal{L}_{3}$ on the graph categories defined in the previous section. For simplicity, we define $\mathcal{L}_{n}^{\prime}$ as the non-trivial action of $\mathcal{L}_{n}$:
@@
$\mathcal{L}_{3}\mathsf{F}_{3,\beta}=(w+3J)\mathsf{F}_{3,\beta},\beta=I,X,XX\ .$ (88)
@@ 4.3 Observables
We write $\mathsf{w}=-\text{i}\mathsf{E}-nJ$, then $M=-\text{i}\mathsf{E}\mathbf{1}+M_{J}$, $M_{J}$ has the explicit form
@@
It is direct to obtain the spectrum of $M_{J}$:
$\Bigg\{$ $0,0,0,-\frac{3(D-2)J}{D},-\frac{3(D-1)J}{D},-\frac{2(D-1)J}{D},-\frac{2(D-1)J}{D},-\frac{2(D-1)J}{D},$ (116)
@@
Then we can find the analytical expression of $\mathcal{U}_{3}(t)$
$e^{\mathcal{L}_{3}t}=e^{-\text{i}\mathsf{E}t}\sum_{a=1}^{26}f_{a}(t)\mathsf{F}_{a}\ .$ (117)
@@
Figure 3: The pattern coefficient $f_{a}(t)$ or its log-plot, where we set $D=10,J=1$. We depict the categories of $p=0,1,2,3$ with red, orange, green and blue color respectively. One can see the three categories with $p=3$ has the energy 0, so they (blue lines) reminds finite at time $t=\infty$.
Generally, we can calculate the observables
$\text{Tr}\left[O_{1}(t)O_{2}O_{3}(t)O_{4}O_{5}(t)O_{6}\right]=\hbox{$\vbox{\hbox{\resizebox{}{}{{\hbox{\set@color{}}}}}}$}\equiv\sum_{a=1}^{26}f_{a}(t)\times\left(e^{-\text{i}\mathsf{E}t}\mathsf{F}_{a}\right)_{O_{1},O_{2},O_{3},O_{4},O_{5},O_{6}}$ (118)
where we have used $O(t)\equiv U^{\dagger}OU$ and the graph representation for an operator AA
$A_{ij}=\hbox{$\vbox{\hbox{\resizebox{}{}{{\hbox{\set@color{}}}}}}$}\ .$ (119)
To calculate $\mathsf{F}_{a}[O_{1},O_{2},O_{3},O_{4},O_{5},O_{6}]$, in principle, we need to contract each individual graph in the category with the operator tensor in the way as shown in Eq.(118). The denotation $\mathsf{F}=\prod_{i=1}^{p}c_{a_{i}\bar{b}_{i}}c[\sigma]$ has told us the connection between points in the graph, we treat the connections on the graph as free path and the operators as bridges which connect points $i,\bar{i}$ and $\bar{i}^{\prime},\overline{(i+1)}^{\prime}$, here prime means the point on the right hand of the graph. So there is no need to draw the graph. We can just start with any points following the connections in the graph then finally we will return back, then we will have an operator trace
@@
If there remind points we do not pass, we then do the same thing until all points in the graph are passed. The process may be a little boring, but it can be achieved by some code. Or one can just write the explicit expression of a graph with Kronecker delta function as defined in Eq.(15).
@@
$\mathbb{E}\left(1_{D}^{\otimes 2}\otimes 1_{D}^{\otimes 2}\otimes U(t_{32})\otimes U^{*}(t_{32})\right).\mathbb{E}\left(1_{D}\otimes 1_{D}\otimes U(t_{21})^{\otimes 2}\otimes U^{*}(t_{21})^{\otimes 2}\right).\mathbb{E}\left(U(t_{1})^{\otimes 3}\otimes U^{*}(t_{1})^{\otimes 3}\right)$
$=$ $\left[1_{D}^{\otimes 2}\otimes 1_{D}^{\otimes 2}\otimes\mathcal{U}_{1}(t_{32})\right].\left[1_{D}\otimes 1_{D}\otimes\mathcal{U}_{2}(t_{21})\right].\mathcal{U}_{3}(t_{1})\ .$ (121)
where $1_{D}$ denotes the $D$-dimensional identity matrix. Finally, we simply need to contract $\mathcal{U}_{3}(t_{1},t_{2},t_{3})$ with the operator tensor as demonstrated in Eq.(118). In this paper, we do not provide explicit expressions for the correlation functions.
5 General case
For general $n$, so we have $2n$ contours $\{1,\bar{1},2,\bar{2},3,\bar{3},\ldots,n,\bar{n}\}$. The permutation group is $S_{n}\otimes S_{n}$. As discussed before, we will meet with terms
$\sum_{u_{1},u_{2},\ldots,u_{p}}c(u_{1},u_{2},\ldots,u_{p})c\left[\prod_{j=1}^{m}(X+\bar{X})^{(\boldsymbol{\alpha}_{j})}\right]$ (122)
where recall $X^{(\boldsymbol{\alpha}_{j})}$ is a set of two-cycles submitting to the condition $\boldsymbol{\alpha}_{j}$, i.e.,